Riverside Inc. makes one model of wooden canoe. Partial information for it follows:

Number of Canoes Produced and Sold
495 645 795
Total costs
Variable costs $ 65,835 ? ?
Fixed costs 149,300 ? ?
Total costs $ 215,135 ? ?
Cost per unit
Variable cost per unit ? ? ?
Fixed cost per unit ? ? ?
Total cost per unit ? ? ?

Required:

1. Complete the table.

3. Suppose Riverside sells its canoes for $507 each. Calculate the contribution margin per canoe and the contribution margin ratio.

4. Next year Riverside expects to sell 845 canoes. Complete the contribution margin income statement for the company.
Answer #1

1.

Number of canoes produced and sold
495 645 795
Total costs:
Variable cost $65,835 $85,785 $105,735
Fixed cost $149,300 $149,300 $149,300
Total cost $215,135 $235,085 $255,035
Variable cost per unit $133 $133 $133
Fixed cost per unit $301.6 $231.47 $187.8
Total cost per unit $434.6 $364.47 $320.8

Working note

1. Variable cost per unit = $65,835 / 495 = $133 , this is constant, doesn't change with number of canoes produced and sold

Variable cost:

for 645 canoes = 645 * 133 = $85,785

for 795 canoes = 795 * 133 = $105,735

2. Fixed cost, $149,300, is constant for 495, 645 and 795 canoes produced and sold. But fixed cost per unit will change according to the number of canoes produced and sold.

Fixed cost per unit:

for 495 canoes = 149,300 / 495 = $301.6

for 645 canoes = 149,300 / 645 = $231.5

for 795 canoes = 149,300 / 795 = 187.8

3.Total cost

495 canoes 645 canoes 795 canoes
Total cost variable cost+fixed cost

68,835+149,300=

215,135

85,785+149,300=

235,085

105,735+149,300=

Total cost per unit variable cost per unit+ fixed cost pe unit 133+301.6=434.6 133+231.47=364.47 133+187.8=320.8

2. If each canoes are sold for $507,

Contribution margin per canoe = selling price - variable cost = 507 - 133 = $374

Contribution margin ratio = Contribution margin / selling price = 374 / 507 = 0.7376

Contribution margin ratio in percentage = 0.7376 * 100 = 73.76%

3. Contribution margin Income statement if sales is 845 canoes

Riverside,inc
Income statement

Sales

($507*845)

$428,415

Less: Variable cost

($133*845)

($112,385)
Contribution margin $316,030
Less: Fixed cost ($149,300)
Net income $166,730